Shift Summary, Owl Nov 15, 1999

Douglas A. Jensen (djensen@fnal.gov)
Mon, 15 Nov 1999 07:04:06 -0600 (CST)

Shift Summary for Midnight-8 shift, 15/Nov/99

On Shift: Rick K, Doug J + Jason L.

Hours of beam available at all intensities: 6.4
Hours of beam available at nominal intensity: 6.4
Hours of E799 data (special runs not included): 4.1
Hours of special runs (laser scans, muon runs...): .5 (looooong mu run)

Accounting of lost time:

Banana, Access to put reg in, take it out, NM3S off

========================================================================
Runs ended this shift:
Run 14973 NM4AN on and sweepers at nominal E799
long muon run on moving regenerator
Run 14974 BEAM_799, some problems

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: two
0030 - Put Regenerator in and out

========================================================================
Overview of shift:

Bananarama under way. Jason L arrives to work on Banana problems.

Make a long muon run with regenerator in and NOMINAL MAGNET SETTINGS.
The reg DEMAND HV is OK, but some of the measured HV readback is zero.
In particular, we are interested in the last few elements;
the measured HV is ok for the bottom but zero for the top,
so this run may [not] be useful for the reg position study.
Since this muon run is special, we just leave it and hope that
the bottom channels are sufficient; if not then the reg can be put
in for later muon runs after fixing the HV problems.
See ~sdaq/hv1440/reg/hv_nov15.txt for MEASURED REG HV.

Also noticed red SDAQ button and lots if missing info like
hall temp and humidty.

The spare ktev2 drive is jammed so we cannot mount the most
recent split tape UPM029. Experts contacted.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14970 1 00:15 End Run. banana problems.
Jason arrives to work on banana problem.

00:30 Access to move regenerator into operating position
Reg HV turnon problems...
00:50 MCR needs to access NM3 to reset rad monitor. waiting
01:05 Have gotten 2 spills, beam stop in for muon running
all magnets nominal
14973 1 01:06 run starts. Tweeking prescales
3 regenerator move staarted

DC4 has tripped off! Page Hogan
TRD 5,6 warning
25 01:37 Angela returns call, Now HV's on DC4 on.

Unplug monitor plane DPL bit because PL3 gets
too much data and causes dead time.
02:15 STOP run with 2M B01 and 2.5M B06 triggers
02:15 Mount UPMU03 and turn TRD HV back on
Call for access to remove REG
02:29 Call to end access, request beam.
new run initialized.
14974 02:30 Start
18 02:55 NM3S starts ramping, MCR called to see if we
wanted it on! Data taking starts here so junk
first 17 spills.
(Suspect that NM3S got turned off when MCR folks
went in to reset Rad Monitor. Didn't get turned
back on...??? BEWARE ! )
53 03:45 KTeVANA errors, please run KFRIEND
75 04:08 getting 'please check L3 accept fraction' messages
so end the run, do a dart fresh start....
14975 0 04:20 Start the new run
reload HCC, a couple of KFRIENDS
6 04:26 Good data starts at spill 6.
04:30 SDAQ process error. checking/waiting
Note that Magnet monitor thinks we are in 14967 !
SDAQ has been 'sleeping on the job' for a while.
Intensity about 8.8E12, BOA 1.09 OK
120 07:00 New Shift crew here, run continues....
========================================================================
Follow up Notes on Previous Shift
Banana problems under investigation.

========================================================================
Items to Follow up on Next Shift
CO2 Dewar looks mighty low. Check, and replace as necessary.
Cosmic ADC plots look to have low amplitudes. time + few nsec.
One channel in RC6 has a high gain. Frown face on RC ADCs
kumquat hardware ineff for 4x a few times more than ref plots.
Hyperon beta trig yield .148, was .223?????